OpenStudy (anonymous):

Evaluate. 8 – 2 • 3 + 7

OpenStudy (anonymous):

8 – 2 • 3 + 7 =8-6+7 =15-6 =9

OpenStudy (anonymous):

ok thanks Learner11

OpenStudy (anonymous):

here is a medal

OpenStudy (anonymous):

no problem . Anytime :)